# refs.bib {refsection}

### v2

@@ 2.3.1 Language Models
We evaluate a diverse collection of state-of-the-art transformer-based LMs (Supplementary Table 5). Our selection criteria included model license, context length, and size. For open-source models, we include Llama touvron2023llama; grattafiori2024llama (Llama 3.1 8B, 3.2 3B, and 3.3 70B), Qwen3 yang2025qwen3 (4B dense version), and Gemma team2025gemma (Gemma 3 27B and MedGemma 27B). For proprietary models, we include GPT-4o Mini, GPT-4o openai2023gpt4; hurst2024gpt, Claude Sonnet 4, and Gemini 2.0 Flash team2023gemini for their enhanced performance capabilities zheng2023judging.
@@ 2.3.2 Implementation Details
Under open-source category, we select LMs under 8 billion parameters for fine-tuning, as they can be easily deployed on consumer-grade GPUs. We leverage DSPy’s bootstrap fine-tuning algorithm khattab2023dspy; soylu2024fine. As part of our contribution, we extend DSPy’s existing local parameter-efficient fine-tuning pipeline to enable Quantized Low-Rank Adaptation (QLoRA) (GitHub PRdettmers2023qlora. We employ an NVIDIA A6000 GPU and fine-tune models for 5 epochs using the Adam optimizer kingma2014adam. We initialize the learning rate at $1\times 10^{-5}$ with linear decay. We set the per-device train batch size to 1 and apply 4-bit precision quantization using BitsAndBytesConfig dettmers2022gpt3.
@@ 2.3.3 Statistical Analysis
We compute the F1 score for each risk level and task independently, comparing LM-predicted risk classifications to physician labels. The F1 score is a measure of a model’s accuracy that considers both precision (the correctness of the positive predictions) and recall (the ability to identify all positive instances). The F1 score, defined as the harmonic mean of precision and recall, ranges from 0 to 1, where 1 indicates perfect precision and recall, and 0 indicates the worst performance with no correct positive predictions. Overall F1 scores are reported using the macro average across tasks, ensuring equal weighting regardless of task size.
@@
We also report Cohen’s $\kappa$, allowing us to assess a LM’s agreement with physicians. Cohen’s $\kappa$ cohen1960coefficient is a statistical measure of agreement for categorical items, which takes into account the possibility of agreement occurring by chance. Given the ordinal nature of the four-class validation task, we compute Cohen’s $\kappa$ with linear weighting, comparing model-predicted risk levels to physician-assigned levels (1–4). The $\kappa$ score ranges from -1 to 1, where 1 represents perfect agreement, 0 indicates agreement no greater than chance, and -1 indicates complete disagreement. We use Krippendorff’s $\alpha$ krippendorff2018content to assess inter-physician agreement over a subset of examples annotated by multiple physicians, evaluating the robustness of the physician annotations. Krippendorff’s $\alpha$ ranges from 0 to 1, where 1 indicates perfect agreement among annotators.
@@
To assess whether MedVAL significantly improves performance over baseline models, we apply the McNemar test, a non-parametric test for paired nominal data that detects differences in classification outcomes. We evaluate whether MedVAL and the baseline model differ in their ability to correctly classify examples. To account for multiple comparisons (models $\times$ risk levels), we apply a Bonferroni correction with a Type I error rate of $\alpha=0.05$. All statistical analyses were conducted using Python’s statsmodels package.
@@
Figure 3: Performance benchmark (F1 score). a) We report the performance of LMs before and after MedVAL fine-tuning. b) We rank all LMs (low to high), grouped into three methods. c) We report the $\Delta$ F1 score between MedVAL fine-tuned and baseline LM performance across each prediction class. Baseline indicates zero-shot LM before fine-tuning, Baseline (larger comparator) indicates a larger zero-shot LM as reference (not chosen for fine-tuning), and MedVAL indicates LM after fine-tuning. ${}^{*}p<0.001$ indicates statistically significant difference in classification performance of MedVAL and baseline (McNemar test). Notably, smaller MedVAL fine-tuned LMs match or exceed the performance of much larger baseline LMs. Furthermore, MedVAL fine-tuned Qwen3-4B (52.7%) and GPT-4o (58.7%) achieve the highest F1 score ranking under respective categories.
@@ 3 Results
We compare two methods: 1) zero-shot baseline, and 2) MedVAL. Baseline involves directly prompting the LM without additional adaptation, whereas MedVAL refers to fine-tuning the model via the MedVAL framework.
@@ 3.1 Overall Performance
We show that MedVAL fine-tuning consistently improves the performance of all LMs. The F1 score and Cohen’s $\kappa$, which reflect classification accuracy and agreement with expert-generated results, respectively, are reported in Figure 3 and Supplementary Figure 5. Across all tested LMs, MedVAL fine-tuning leads to average baseline F1 scores improving from 36.7% to 51.0%. On average, MedVAL improved F1 scores by 113% for open-source LMs and 11% for proprietary LMs, where baseline performance was already considerably high. Despite the inherent difficulty of this four-class classification task, where even frontier LMs struggle and baseline scores never exceed 55%, MedVAL fine-tuning improves GPT-4o’s F1 score from 54.5% to 58.7%, establishing the best performance in our benchmark.
First, we examine performance across open-source models. Among all open-source LMs, Qwen-3-4B exhibits the highest F1 score after MedVAL fine-tuning (42.8% $\to$ 52.7%), surpassing other notably larger open-source models such as Llama-3.3-70B and even the medically fine-tuned MedGemma-27B. Among zero-shot baselines, MedGemma-27B achieves the highest performance (48.2%), slightly outperforming Llama-3.3-70B and Gemma-3-27B. Notably, MedVAL fine-tuning substantially improves the F1 scores of Llama-3.2-3B (12.8% $\to$ 42.9%), Llama-3.1-8B (25.9% $\to$ 46.5%). Despite Llama-3.1-8B also benefiting from MedVAL, MedVAL Qwen-3-4B outperforms it, emphasizing the advantage of specific baseline LM training techniques. Remarkably, MedVAL Qwen-3-4B even outshines certain proprietary baselines such as GPT-4o Mini (47.4%) and Gemini 2.0 Flash (51.5%). We refer to the MedVAL fine-tuned Qwen-3-4B model as “MedVAL-4B” for public release.
Within proprietary LMs, post MedVAL fine-tuning, GPT-4o Mini (47.4% $\to$ 54.0%) rivals the baseline performance of larger Claude Sonnet 4 (54.2%) and GPT-4o (54.5%), and surpasses Gemini 2.0 Flash (51.5%), highlighting that MedVAL can enable LMs to reach or exceed baseline performance of much larger LMs.
We report Krippendorff’s $\alpha$ to assess inter-reader variability across the six tasks in Table 3a. The overall $\alpha$ of 0.848 indicates substantial to almost perfect agreement, proving the reliability of the physician annotations. A value of $\alpha\geq 0.80$ indicates reliable rating quality to draw triangulated conclusions based on the rated data marzi2024k. Notably, the highest inter-physician agreement was observed in the bhc2spanish and medication2answer tasks, with an $\alpha$ of 0.943 and 0.904, respectively, while the lowest was in the query2question task, with an $\alpha$ of 0.560. We also observe that physician agreement correlates strongly with MedVAL performance: we observe a Pearson correlation of $r=0.67$ between inter-physician agreement and GPT-4o MedVAL F1 scores. This suggests that when experts agree more, MedVAL produces higher and consistent F1 scores.
@@ 3.2 Risk-Level Classification Performance
Proprietary models like GPT-4o Mini exhibited notable gains at intermediate levels (level 2: 14.5%; level 3: 13.2%). Even GPT-4o, the strongest baseline model, showed smaller but consistent improvements, especially at level 2 (7.3%) and level 3 (6.9%). While these may appear modest, they are clinically significant: levels 2 and 3 represent the threshold between outputs that are likely safe for use with minimal oversight (level 2) and those requiring human review or revision (level 3). Thus, MedVAL enhances model sensitivity to this crucial safety boundary, improving decision-making precision at a critical deployment threshold. Although minor drops were observed in a few model-risk combinations, such as GPT-4o Mini at level 4 (-2.2%) and Llama-3.2-3B at level 3 (-2.7%), the overall pattern shows strong and consistent gains across clinically meaningful risk categories.
Table 3: a) Task-wise performance benchmark (F1 score). Bolded values highlight best performance under respective categories (open-source/proprietary). Green and red values indicate improvement or decline in F1 score compared to the corresponding baseline LM. On average across LMs, MedVAL fine-tuning increases F1 scores by 65% on in-distribution tasks and 84% on out-of-distribution tasks. Krippendorff’s $\alpha=0.848$ represents high inter-physician agreement across examples evaluated by multiple physicians, indicating the reliability of physician annotations. MedVAL fine-tuned Qwen3-4B (referred to as MedVAL-4B) achieves the best performance among all open-source LMs, matching the baseline performance of proprietary LMs.
@@
# Model Method Sensitivity Specificity F1 Score Accuracy
1 Llama-3.2-3B Baseline 0.086$\pm$0.01 0.960$\pm$0.01 0.153$\pm$0.02 0.474$\pm$0.02
MedVAL 0.919$\pm$0.01 0.560$\pm$0.02 0.809$\pm$0.01 0.760$\pm$0.01
2 Llama-3.1-8B Baseline 0.670$\pm$0.02 0.651$\pm$0.03 0.688$\pm$0.02 0.662$\pm$0.02
MedVAL 0.788$\pm$0.02 0.786$\pm$0.01 0.804$\pm$0.01 0.787$\pm$0.01
3 Qwen3-4B Baseline 0.858$\pm$0.02 0.643$\pm$0.03 0.800$\pm$0.01 0.762$\pm$0.02
MedVAL 0.839$\pm$0.02 0.752$\pm$0.02 0.823$\pm$0.01 0.800$\pm$0.01
Ensemble (1+2+3) MedVAL 0.899$\pm$0.02 0.686$\pm$0.03 0.837$\pm$0.01 0.805$\pm$0.01
4 GPT-4o Mini Baseline 0.784$\pm$0.02 0.807$\pm$0.02 0.809$\pm$0.02 0.794$\pm$0.02
MedVAL 0.848$\pm$0.02 0.831$\pm$0.02 0.855$\pm$0.01 0.840$\pm$0.01
5 GPT-4o Baseline 0.835$\pm$0.02 0.861$\pm$0.02 0.858$\pm$0.01 0.846$\pm$0.01
MedVAL 0.792$\pm$0.02 0.906$\pm$0.02 0.849$\pm$0.01 0.843$\pm$0.01
Ensemble (4+5) MedVAL 0.874$\pm$0.02 0.815$\pm$0.02 0.864$\pm$0.01 0.848$\pm$0.01
@@ 3.3 Task-Wise Performance
The task-wise F1 scores illustrated in Table 3a (standard deviations in Supplementary Table 6) reveal substantial improvement in LM performance across tasks. On average across LMs, MedVAL fine-tuning yielded a 65% F1 score improvement on in-distribution tasks and an 84% improvement on out-of-distribution tasks. Notably, MedVAL displays strong improvements on dialogue2note, an unseen task with the longest input context lengths (average $\sim 1.5$k tokens), underscoring robustness in highly complex tasks.
@@
In the proprietary category, MedVAL GPT-4o saw a consistent performance increase across all tasks, leading to the highest overall F1 score (58.7%), even though the LM had the best performance in only 2 out of 6 tasks (medication2answer and impression2simplified). Notably, GPT-4o Mini outperformed all LMs on two tasks (report2impression and dialogue2note) after MedVAL fine-tuning. While GPT-4o-mini saw performance gains across tasks after MedVAL, it also saw minor drops in query2question and bhc2spanish. However, we observed the query2question task showed consistently lower performance across any model, aligning with its low inter-annotator agreement ($\alpha=0.560$).
We also conducted an in-depth analysis to understand the implications of LM predictions. For a granular understanding, we reviewed examples where the baseline LM incorrectly evaluated the output, but corrected the evaluation after MedVAL fine-tuning. We present such examples in Figure 4 and Supplementary Figure 7. Together, these findings demonstrate that MedVAL enhances both the generalization and task-specific capabilities of diverse LMs, bringing even modestly-sized LMs closer to expert-level accuracy.
Figure 4: Representative examples of validation of LM-generated medical text by 1) the physician, 2) baseline GPT-4o, and 3) MedVAL fine-tuned GPT-4o. Under each example, MedVAL demonstrates higher agreement with the physician.
@@ 3.4 Safety (Binary) Classification Performance
We assess how well LMs can distinguish between safe (risk levels 1–2) and unsafe (levels 3–4) outputs in a binary deployment setting. As shown in Table 3b, MedVAL fine-tuning consistently improves binary classification performance across most LMs. Across all the LMs fine-tuned, MedVAL improved average accuracy from 70.8% to 80.6% and F1 score from 66.2% to 82.8%, demonstrating more reliable discrimination between deployment-safe and unsafe responses. Notably, ensembling MedVAL LM outputs across all open-source models (rows 1–3) and, separately, across all proprietary models (rows 4–5) yielded the highest F1/accuracy within each category (83.7% / 80.5% for open-source and 86.4% / 84.8% for proprietary).
Among open-source LMs, the strongest gains were seen in Llama-3.2-3B (F1 score 15.3% $\to$ 80.9%). Similarly, F1 scores of Llama-3.1-8B improved from 68.8% to 80.4%, and Qwen3-4B from 80.0% to 82.3%. For proprietary models, GPT-4o Mini showed steady gains in all metrics, rising from 80.9% to 85.5% in F1 score. GPT-4o, which already had high baseline performance (F1: 85.8%, accuracy: 84.6%), retained strong performance with only marginal changes (F1: 84.9%, accuracy: 84.3%). Overall, we show that MedVAL translates to more accurate safety-critical deployment decisions at a per-sample level.

### v4

@@ 2.3.1 Language Models
We evaluate a diverse collection of state-of-the-art transformer-based LMs (Supplementary Table 5). Our selection criteria included model license, context length, and size. For open-source models, we include Llama [55, 56] (Llama 3.1 8B, 3.2 3B, and 3.3 70B), Qwen3 [57] (4B dense version), and Gemma [58] (Gemma 3 27B and MedGemma 27B). For proprietary models, we include GPT-4o Mini, GPT-4o [59, 60], Claude Sonnet 4, and Gemini 2.0 Flash [61] for their enhanced performance capabilities [62].
@@ 2.3.2 Implementation Details
Under open-source category, we select LMs under 8 billion parameters for fine-tuning, as they can be easily deployed on consumer-grade GPUs. We leverage DSPy’s bootstrap fine-tuning algorithm [63, 64]. As part of our contribution, we extend DSPy’s existing local parameter-efficient fine-tuning pipeline to enable Quantized Low-Rank Adaptation (QLoRA) (GitHub PR) [65]. We employ an NVIDIA A6000 GPU and fine-tune models for 5 epochs using the Adam optimizer [66]. We initialize the learning rate at $1\times 10^{-5}$ with linear decay. We set the per-device train batch size to 1 and apply 4-bit precision quantization using BitsAndBytesConfig [67].
@@ 2.3.3 Statistical Analysis
We compute the F1 score for each risk level and task independently, comparing LM-predicted risk classifications to physician labels. The F1 score is a measure of a model’s accuracy that considers both precision (the correctness of the positive predictions) and recall (the ability to identify all positive instances). F1 score ranges from 0 to 1, where 1 indicates perfect precision and recall, and 0 indicates the worst performance with no correct positive predictions. We report F1 score using the macro average, ensuring equal weighting for risk levels and tasks to help mitigate class imbalance. For evaluating binary safety classification, we additionally report sensitivity (unsafe recall) and specificity (safe recall) to make class balance explicit.
@@
We also report Cohen’s $\kappa$, allowing us to assess a LM’s agreement with physicians. Cohen’s $\kappa$ [68] is a statistical measure of agreement for categorical items, which takes into account the possibility of agreement occurring by chance. Given the ordinal nature of the four-class validation task, we compute Cohen’s $\kappa$ with linear weighting, comparing model-predicted risk levels to physician-assigned levels (1–4). The $\kappa$ score ranges from -1 to 1, where 1 represents perfect agreement, 0 indicates agreement no greater than chance, and -1 indicates complete disagreement. We use Krippendorff’s $\alpha$ [69] to assess inter-physician agreement over a subset of examples annotated by multiple physicians, evaluating the robustness of the physician annotations. Krippendorff’s $\alpha$ ranges from 0 to 1, where 1 indicates perfect agreement among annotators.
@@
To assess whether MedVAL significantly improves performance over baseline models, we apply the McNemar test, a non-parametric test for paired nominal data that detects differences in classification outcomes. We evaluate whether MedVAL and the baseline model differ in their ability to correctly classify examples. To account for multiple comparisons (10 models $\times$ 4 risk levels), we apply a Bonferroni correction with a Type I error rate of $\alpha=0.05$. All statistical analyses were conducted using Python’s statsmodels package.
To test the non-inferiority (NI) of a LM against a single human expert, we use the 90/840 test cases with multiple-physician ratings. To identify a reference label, we use majority consensus. A "single expert" is defined by randomly sampling one physician’s label. We compute macro-F1 against the reference for both the model and sampled expert, form the NI contrast $\Delta$ = F1 (model) - F1 (single expert), and use a paired item-level bootstrap (B=10,000). Non-inferiority is declared if $\text{LCB}_{95\%}(\Delta)>-0.05$ (clinically negligible difference). We did not apply a multiplicity adjustment as our NI comparison was GPT-4o vs a single expert.
@@
Figure 3: Performance benchmark (F1 score). a) We report the performance of LMs before and after MedVAL distillation. b) We rank all LMs (low to high), grouped into three methods. c) We report the $\Delta$ F1 score between MedVAL and baseline LM performance across each prediction class. Baseline indicates zero-shot LM before distillation, Baseline (larger comparator) indicates a larger zero-shot LM as reference (not chosen for distillation), and MedVAL indicates LM after distillation. $^{*}p<0.001$ indicates statistically significant difference in classification performance of MedVAL and baseline (McNemar test). Notably, smaller MedVAL LMs match or exceed the performance of much larger baseline LMs. Furthermore, MedVAL Qwen3-4B (52.7%) and GPT-4o (58.7%) achieve the highest F1 score ranking under respective categories.
@@ 3 Results
We compare two methods: 1) zero-shot baseline, and 2) MedVAL. Baseline involves directly prompting the LM without additional adaptation, whereas MedVAL refers to fine-tuning the model via MedVAL distillation.
@@ 3.1 Overall Performance
We show that MedVAL distillation consistently improves the performance of all LMs. The F1 score and Cohen’s $\kappa$, which reflect classification accuracy and agreement with expert-generated results, respectively, are reported in Figure 3 and Supplementary Figure 5. Across all tested LMs, MedVAL distillation leads to average baseline F1 scores improving from 36.7% to 51.0%. On average, MedVAL improved F1 scores by 113% for open-source LMs and 11% for proprietary LMs, where baseline performance was already considerably high. Despite the inherent difficulty of this four-class classification task, where even frontier LMs struggle and baseline scores never exceed 55%, MedVAL distillation improves GPT-4o’s F1 score from 54.5% to 58.7%, establishing the best performance in our benchmark.
First, we examine performance across open-source models. Among all open-source LMs, Qwen-3-4B exhibits the highest F1 score after MedVAL distillation (42.8% $\to$ 52.7%), surpassing other notably larger open-source models such as Llama-3.3-70B and even the medically fine-tuned MedGemma-27B. Among zero-shot baselines, MedGemma-27B achieves the highest performance (48.2%), slightly outperforming Llama-3.3-70B and Gemma-3-27B. Notably, MedVAL distillation substantially improves the F1 scores of Llama-3.2-3B (12.8% $\to$ 42.9%), Llama-3.1-8B (25.9% $\to$ 46.5%). Despite Llama-3.1-8B also benefiting from MedVAL, MedVAL Qwen-3-4B outperforms it, emphasizing the advantage of specific baseline LM training techniques. Remarkably, MedVAL Qwen-3-4B even outshines certain proprietary baselines such as GPT-4o Mini (47.4%) and Gemini 2.0 Flash (51.5%). We refer to the MedVAL Qwen-3-4B model as “MedVAL-4B” for public release.
Within proprietary LMs, post MedVAL distillation, GPT-4o Mini (47.4% $\to$ 54.0%) rivals the baseline performance of larger Claude Sonnet 4 (54.2%) and GPT-4o (54.5%), and surpasses Gemini 2.0 Flash (51.5%), highlighting that MedVAL can enable LMs to reach or exceed baseline performance of much larger LMs.
We report Krippendorff’s $\alpha$ to assess inter-reader variability across the six tasks in Table 3a. The overall $\alpha$ of 0.848 indicates substantial to almost perfect agreement, proving the reliability of the physician annotations. A value of $\alpha\geq 0.80$ indicates reliable rating quality to draw triangulated conclusions based on the rated data [70]. Notably, the highest inter-physician agreement was observed in the bhc2spanish and medication2answer tasks, with an $\alpha$ of 0.943 and 0.904, respectively, while the lowest was in the query2question task, with an $\alpha$ of 0.560. We also observe that physician agreement correlates strongly with MedVAL performance: we observe a Pearson correlation of $r=0.67$ between inter-physician agreement and GPT-4o MedVAL F1 scores. This suggests that when experts agree more, MedVAL produces higher and consistent F1 scores.
@@ 3.2 Risk-Level Classification Performance
Proprietary models like GPT-4o Mini exhibited notable gains at intermediate levels (level 2: 14.5%; level 3: 13.2%). Even GPT-4o, the strongest baseline model, showed smaller but consistent improvements, especially at level 2 (7.3%) and level 3 (6.9%). These gains are clinically significant because levels 2 and 3 represent the threshold between outputs that are likely safe for use with minimal oversight (level 2) and those requiring human review or revision (level 3). Interestingly, MedVAL sharpens this decision boundary for even high-performing baseline frontier LMs. Although minor drops were observed in a few model-risk combinations, such as GPT-4o Mini at level 4 (-2.2%) and Llama-3.2-3B at level 3 (-2.7%), the overall pattern shows strong and consistent gains across clinically meaningful risk categories.
Table 3: a) Task-wise performance benchmark (F1 score). Bolded values highlight best performance under respective categories (open-source/proprietary). Green and red values indicate improvement or decline in F1 score compared to the corresponding baseline LM. On average across LMs, MedVAL distillation increases F1 scores by 65% on in-distribution tasks and 84% on out-of-distribution tasks. Krippendorff’s $\alpha=0.848$ represents high inter-physician agreement across examples evaluated by multiple physicians, indicating the reliability of physician annotations. MedVAL Qwen3-4B (referred to as MedVAL-4B) achieves the best performance among all open-source LMs, matching the baseline performance of proprietary LMs.
@@
b) Safe/unsafe (binary) performance benchmark. Bolded values represent the best-performing method under each model (MedVAL vs Baseline). Green values represent the best-performing model/method combination under respective categories (open-source/proprietary). Ensemble indicates aggregation of outputs from multiple MedVAL LMs. Across all LMs, MedVAL improves average accuracy from 71% to 81% and F1 score from 66% to 83%, demonstrating reliable discrimination between deployment-safe and unsafe responses. Notably, inter-physician agreement (Krippendorff’s $\alpha=0.754$) falls within range of the best MedVAL classifiers (Accuracy/F1 $80\%)$, suggesting comparable consistency, albeit measured via a different metric.
# Model Method Sensitivity Specificity F1 Score Accuracy
@@
[…]
@@ 3.3 Task-Wise Performance
The task-wise F1 scores illustrated in Table 3a (standard deviations in Supplementary Table 7) reveal substantial improvement in LM performance. On average across LMs, MedVAL distillation yielded a 65% F1 score improvement on seen tasks and an 84% improvement on unseen tasks. Notably, MedVAL displays strong improvements on dialogue2note, an unseen task with the longest input context lengths (average $\sim 1.5$k).
@@
In the proprietary category, MedVAL GPT-4o saw a consistent performance increase across all tasks, leading to the highest overall F1 score (58.7%). Notably, MedVAL GPT-4o Mini outperforms all LMs on two tasks (report2impression and dialogue2note). However, MedVAL GPT-4o Mini also sees minor drops in query2question and bhc2spanish. However, we observe that the query2question task shows consistently lower performance across models, aligning with its low inter-annotator agreement ($\alpha=0.560$).
We also conduct an in-depth analysis to understand the implications of LM predictions. For a granular understanding, we review examples where the baseline LM incorrectly evaluates the output, but corrects the evaluation after MedVAL distillation. We present such examples in Figure 4, Supplementary Figure 7, and Supplementary Figure 8 (showing "reasoning" behind LM’s risk-grade prediction). Together, these findings demonstrate that MedVAL enhances both the generalization and task-specific capabilities of diverse LMs.
Figure 4: Representative examples of validation of LM-generated medical text by 1) the physician, 2) baseline GPT-4o, and 3) MedVAL GPT-4o. Under each example, MedVAL demonstrates higher agreement with the physician.
@@ 3.4 Safety (Binary) Classification Performance
We assess how well LMs can distinguish between safe (risk levels 1–2) and unsafe (levels 3–4) outputs in a binary deployment setting. As shown in Table 3b, MedVAL distillation consistently improves binary classification performance across most LMs. MedVAL improves average accuracy from 70.8% to 80.6% and F1 score from 66.2% to 82.8%, demonstrating more reliable discrimination between safe and unsafe responses. Notably, ensembling MedVAL LM outputs across all open-source models (rows 1–3) and, separately, across all proprietary models (rows 4–5) yields the highest F1/accuracy within each category.
Among open-source LMs, the strongest gains are seen in Llama-3.2-3B (F1 score 15.3% $\to$ 80.9%). Similarly, F1 scores of Llama-3.1-8B improve from 68.8% to 80.4%, and Qwen3-4B from 80.0% to 82.3%. For proprietary models, GPT-4o Mini shows steady gains in all metrics, rising from 80.9% to 85.5% in F1 score. GPT-4o, which already has high baseline performance (F1: 85.8%, accuracy: 84.6%), demonstrates only marginal changes (F1: 84.9%, accuracy: 84.3%). Overall, we show that MedVAL enables more accurate deployment decisions at a per-sample level.